Sam is 9 years older than nick. 7 years ago he was 4 times as old as nick. how old is same now?
Answer by algebrahouse.com(1659)

x = Nick's age now
x + 9 = Sam's age now {Sam is 9 years older than Nick}


x - 7 = Nick's age 7 years ago
x + 2 = Sam's age 7 years ago {subtracted 7 from age now}

x + 2 = 4(x - 7) {seven years ago, Sam was four times Nick}
x + 2 = 4x - 28 {used distributive property}
3x = 30 {added 28 and subtracted x from each side}
x = 10 {divided each side by 3}
x + 9 = 19 {substituted 10, in for x, into x + 9}

Sam is 19

For more help from me, visit: www.algebrahouse.com